You have assigned the following values to these three firms:

Price Upcoming Dividend Growth Beta
US Bancorp $ 29.80 $ 3.20 8.80 % 1.59
Praxair 59.15 1.51 13.00 2.20
Eastman Kodak 36.50 1.00 11.50 0.98

Assume that the market portfolio will earn 15.50 percent and the risk-free rate is 5.70 percent.

Compute the required return for each company using both CAPM and the constant-growth model. (Do not round intermediate calculations and round your final answers to 2 decimal places.)  
Answer #1

CAPM:
1.
=5.70%+1.59*(15.50%-5.70%)
=21.28%

2.
=5.70%+2.20*(15.50%-5.70%)
=27.26%

3.
=5.70%+0.98*(15.50%-5.70%)
=15.30%

Constant growth model:
1.
=3.20/29.80+8.80%
=19.54%

2.
=1.51/59.15+13%
=15.55%

3.
=1.00/36.50+11.50%
=14.24%